# New Coevolution Dynamic as an Optimization Strategy in Group Problem Solving

### v1

@@ 3 Results
  • Reference Case (RC): dynamics with no coevolution, equivalent to $L=0$.
The model was implemented using Fortran and the analysis and graphics were developed using Python (packages: Numpy [21], Matplotlib [22], NetworkX [23] and Pandas [24]) and Gnuplot. All codes and data are available upon request.
@@ 3.1 Reference case
Firstly we present the results with no coevolution: the Reference Case (RC). At $K=0$ (Figure 2(a)) the best result (lower $\left\langle C\right\rangle$) is given by the full connected network (FCN) and the worst by the independent search (IS). The increasing average degree $\mu$ interpolates from IS to the FCN case. As there is only one maximum (the global one) it becomes easier to increase the fitness with more neighbors to copy. However, in terms of system size $M$ there is one optimal value roughly between 15 and 20 which returns $\left\langle C\right\rangle\sim 0.007$. So, if there is too few neighbors (smaller $M$) or too much ones (larger $M$) the agents spend more time looking for the global maximum. At large $M\gtrsim 10^{4}$ the system behaviors as the limit case $M/2^{N}$.
Figure 2: Computational cost $\left\langle C\right\rangle$ as function of $M$ for different average degrees $\mu$ of the ERN with no coevolution. Parameter: $q=1$ and $S=5\times 10^{4}$ samples. The green line is the IS (Eq. 1), black dashed line is the FCN result and the pink line is the approximation for large $M$. Purple, blue, green, orange and red circles correspond to an average degree of $\mu=1$, 5, 10, 20 and 100, respectively. (a) $K=0$. (b) $K=4$.
When $K=4$ (Figure 2(b)) the scenario is drastically different. As before there is an optimal value of $M$ roughly between 10 and 15. However there is a significant increase on the cost $\left\langle C\right\rangle$ at $M\gtrsim 20$ when $\mu$ is increased towards the FCN. As there are local maxima in this landscape, an agent can be trapped in one of these making it longer to find the global maxima. The results on these two scenario resemble the similar case on the random geometric graph [25].
@@ 3.2 Dynamics with coevolution
In order to better enhance the effect of coevolution on the cost, we restricted the values of the average degree for the analysis: we have used $\mu=1.0$ and $\mu=10.0$ for the cases $K=0$ and $K=4$, respectively. We begin our study with the linear case: $q=1$. Considering the easy landscape with $K=0$ (Fig. 3(a)), the rewiring brings benefit to the cost except when $L=1$. As there is only one maximum, any other agent with a higher fitness is worth copying. And this process is facilitated by the selective rewiring of the coevolution. Interestingly, the case $L=1$ increases the cost. The difficult case $K=4$ (Fig. 3(b)) has some similarities: the cost increases at $L=1$ and then decreases up to $L=5$. At $L=6$ there is a significant increase only at $M\sim 60$.
Figure 3: Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of $L$. $L=0$ means no coevolution. Parameters: $q=1$, $S=10^{5}$ samples and $\mu=1$. The green line is the IS (Eq. 1) and the black dashed line is the FCN result. (a) $K=0$. (b) $K=4$.
At higher average degree $\mu$ the results are drastically different. The ease case $K=0$ (Fig. 4(a)) shows influence of $L$ only on small network sizes around $M\lesssim 15$. At $K=4$ the results become more interesting. At small networks coevolution continues to present a higher computational cost. However, at medium-sized networks ($20\leq M\leq 120$), we see a significant improvement in computational cost. $\left\langle C\right\rangle$ presents a maximum at $M=67$ for $L=0$ (Fig. 4(b)) and a minimum at $L=3$.
Figure 4: Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of the number of modifications $L$. Parameters: and $S=10^{5}$ and $\mu=10$. The green line is the IS (Eq. 1) and the black dashed line is the FCN result. (a) $K=0$. (b) $K=4$.
The $\left\langle C\right\rangle$ minimum created by $L$ has a larger sensitivity with $M$ than with $\mu$. As can be seen on Fig. 5(a), the cost has a pronounced minimum between $L=2$ and 3 considering different values of $M$. However when $M$ if fixed and $\mu$ is varied the minimum is not so evident (Fig. 5(b)), and even not present at some values.
Figure 5: Results of $\left\langle C\right\rangle$ vs. $L$. Parameters: $K=4$, $q=0$ and $S=30000$. (a) Different values of $M$ and $\mu=10$. (b) Different values of $\mu$ and $M=67$.
To better visualize the behavior of the $\left\langle C\right\rangle$ minimum at $M=67$, Fig. 6(a) displays a color map on the space $L,\mu$. However, a more interesting measure is the difference between the $L=3$ and $L=0$ values, defined as:
@@
This quantity gives exactly the improvement (given by $\Delta<0$) on the cost $\left\langle C\right\rangle$ created by the coevolution, as can be seen on Fig. 6(b). The average degree as a great impact on this gain: at $L=6$ it goes from destroying at $\mu=6$ to improving the performance at $\mu=12$. This result shows the tunable coevolution described here is a real tool to increase perfomance on the NK model.
Figure 6: Results of $\left\langle C\right\rangle$ in the $(L,\mu)$ space. Parameters: $K=4$, $q=1$, $M=67$ and $S=10^{5}$. (a) $\log\left\langle C\right\rangle$. (b) $\Delta$ defined on Eq. 3. Positive values are indicated in red and negative values in blue.
@@ 3.3 Other dynamics
To verify the impact of the coevolution dynamics, we compare it with the constant rewiring one: $q=0$. Considering the curve $L=1$ from Fig. 4(b), the constant rewiring case gives a lower cost, as can be seen on Fig. 7(a). Focusing on the peak at $M\sim 67$, the results shows a slightly better performance (lower cost) with $L=1$ for the constant rewiring case, while negligible differences when $L=3$. Figures 7(b) and 7(c) show these same results for different average degrees $\mu$. One can see that the improvement from coevolution (or the constant rewiring) case is not present.
Figure 7: Computational cost ($\left\langle C\right\rangle$) as a function of network size $M$ for different values of the average degree. The green line is the IS (Eq. 1), the black dashed line is the FCN result and RC is the reference case (equivalent to $L=0$). In (a) $\mu=10$, (b) $\mu=5$ and (c) $\mu=15$.
Lastly we studied the impact of non-linearity on the computational cost, which happens when $q$ is different from 0 or 1. In this work non-linearity yield no significant outcome, as can be shown on Fig. 8. The case $q=0.5$ only showed slightly better performance than the linear case $q=1$ around $M=67$. Other than this region there were no significant differences.
Figure 8: Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of the non-linearity parameter $q$. Parameters: $\mu=10$, $S=3000$ and $L=3$. The violet circles correspond to $q=0.5$, the orange circles to $q=1.5$, and the red circles to $q=2$. The reference case RC ($L=0$ and $q=1$) is depicted in gray.
@@ 4 Conclusions
In this work, we investigate the impact of coevolution on the NK model within the context of an Erdös-Rényi network. Given that the NK model operates as a fitness model, our goal was to evaluate whether a tunable coevolution yields any gain in computational cost $C$. To this end, we introduce a new dynamic that incorporates the effect of rewiring by changing the neighborhood of the target agent. Notably, this study represents the first exploration of coevolution within the NK model using an ER network, along with the introduction of a new parameter linked to rewiring: the tunable parameter $L$. This parameter allows for rewiring interactions not limited to individual pairs of agents. Our simulations showed that the proposed dynamics can return a reduction in cost on the system. We also investigated the scenario with constant rewiring and have observed that the results are similar to the coevolution case. More studies can be needed to map the regions with improved perfomance for larger $L$. It can be also interesting to study the impact of this tunable coevolution on the phase diagram of with absorbing phase models.
@@ Acknowledgements
The model was implemented using Fortran and the analysis and graphics were developed using Python (packages: Numpy [21], Matplotlib [22], NetworkX [23] and Pandas [24]) and Gnuplot. All codes and data are available upon request.

### v2

@@ 3.1 Benchmark
  • Reference Case (RC): dynamics with a static network, no rewiring, equivalent to $L=0$. Coevolution will bring an advantage if it entails a lower computational cost $C$ than the Reference Case one.
The independent search is the simplest method, and the cheapest one if we think in terms of a company hiring employees to perform the task. The FCN case is also a simple one, however not the cheapest one. As $M$ increases, the cost to enable everybody to communicate with everybody also increases. Hence, for any particular dynamics to be called efficient, it must bring more efficiency than these two cases.
@@ 3.2 Reference case
Firstly we present the results with a static network: the Reference Case (RC). It is important to understand this case so we can highlight the effect of the coevolution. At $K=0$ (Figure 2(a)) the best result (lower $\left\langle C\right\rangle$) is given by the full connected network (FCN) and the worst by the independent search (IS). The increasing average degree $\mu$ interpolates from IS to the FCN case. As there is only one maximum (the global one) it becomes easier to increase the fitness with more neighbors to copy. However, in terms of system size $M$ there is one optimal value roughly between 15 and 20 which returns $\left\langle C\right\rangle\sim 0.007$. So, if there is too few neighbors (smaller $M$) or too much neighbors (larger $M$) the agents spend more time looking for the global maximum. At large $M\gtrsim 10^{4}$ the system behaviors as the limit case $M/2^{N}$.
Figure 4: Coevolution results in a low degree network ($\mu=1$) on the two investigated landscapes. Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of modification $L$, limited by $L<k_{i}$ for each agent. $L=0$ means no coevolution. Parameters: $S=10^{5}$ samples and $\mu=1$. The green line is the IS (Eq. 4) and the black dashed line is the FCN result. (a) Easy landscape: $K=0$. (b) Difficult landscape: $K=4$.
When $K=4$ (Figure 2(b)) the scenario is drastically different. As before there is an optimal value of $M$ roughly between 10 and 15. However there is a significant increase on the cost $\left\langle C\right\rangle$ at $M\gtrsim 20$ when $\mu$ is increased towards the FCN. As there are local maxima in this landscape, an agent can be trapped in one of these making it longer to find the global maxima. The results on these two scenarios resemble the similar case on the random geometric graph [35].
@@ 3.3 Dynamics with coevolution
To enhance the effect of coevolution, we restricted the values of the average degree for the analysis. We observed different scenarios for different values of this degree: coevolution works for some cases and does not work at other cases. We say it works when the computational cost $\left\langle C\right\rangle$ is lower compared to the Reference Case, keeping all other parameters the same. We found local improvements on $\left\langle C\right\rangle$ at the degrees $\mu=1.0$ and $\mu=10.0$ (see Fig. 3). These are the scenarios where coevolution works (at some values of $M$), depicted in Figs. 4, 5 and 8(a). The scenarios where it does not work are depicted in Figs. 8(b) and 8(c).
Let’s start with the low average degree scenario, Fig. 3(a). Considering the easy landscape with $K=0$ (Fig. 4(a)), the coevolution brings benefit to the cost except when $L=1$. As there is only one maximum, any other agent with a higher fitness is worth copying. And this process is facilitated by the selective rewiring of the coevolution. Interestingly, the case $L=1$ increases the cost. The difficult case $K=4$ (Fig. 4(b)) has some similarities: the cost increases at $L=1$ and then decreases up to $L=5$. At $L=6$ there is a significant increase only at $M\sim 60$.
At higher average degree (Fig. 3(b)) the results are drastically different. The ease case $K=0$ (Fig. 5(a)) shows influence of $L$ only on small network sizes around $M\lesssim 15$. At $K=4$ the results become more interesting. At small networks, coevolution continues to present a higher computational cost. However, at medium-sized networks ($20\leq M\leq 120$), we see a significant improvement in computational cost: $\left\langle C\right\rangle$ presents a maximum at $M=67$ for $L=0$ (Fig. 5(b)) and a minimum at $L=3$.
Figure 5: Coevolution results in a high degree network ($\mu=10$) on the two investigated landscapes. Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of modification $L$, limited by $L<k_{i}$ for each agent. $L=0$ means no coevolution. Parameters: $S=10^{5}$ and $\mu=10$. The green line is the IS (Eq. 4) and the black dashed line is the FCN result. (a) Easy landscape: $K=0$. (b) Difficult landscape: $K=4$.
The minimum of $\left\langle C\right\rangle$ with respect to $L$ has a larger sensitivity to $M$ than with to $\mu$. As can be seen on Fig. 6(a), the cost has a pronounced minimum between $L=2$ and 3 considering different values of $M$. However, when $M$ is fixed and $\mu$ is varied, the minimum is not so evident (Fig. 6(b)), and even not present at some values.
Figure 6: Characterization of the computational cost minimum caused by the coevolution. Computational cost $\left\langle C\right\rangle$ vs. $L$. Parameters: $K=4$, $q=0$ and $S=30000$. (a) Minimum on the $(M,L)$ space. Different values of $M$ while $\mu=10$. (b) Different values of $\mu$ while $M=67$.
To better visualize the behavior of the $\left\langle C\right\rangle$ minimum at $M=67$, Fig. 7(a) displays a color map on the space $(L,\mu)$. However, a more interesting measure is the difference between the $L=3$ and $L=0$ values, defined as:
@@
This quantity gives the improvement ($\Delta<0$) on the cost $\left\langle C\right\rangle$ created by the coevolution, as can be seen on Fig. 7(b). The average degree has a great impact on this gain: at $L=6$ it goes from destroying it at $\mu=6$ to improving the performance at $\mu=12$. This result shows the tunable coevolution described here is a real tool to increase perfomance on the NK model.
Figure 7: 2D map of the computational cost $\left\langle C\right\rangle$ and its gain $\Delta$ in the $(L,\mu)$ space. Parameters: $K=4$, $q=1$, $M=67$ and $S=10^{5}$. (a) $\log\left\langle C\right\rangle$. (b) $\Delta$ defined on Eq. 8. Positive values are indicated in red and negative values in blue.
@@ 3.4 Other dynamics
We now verify the differences between the coevolution ($q=1$) and constant rewiring ($q=0$) dynamics. Considering the curve $L=1$ from Fig. 5(b), the constant rewiring case gives a lower cost, as can be seen on Fig. 8(a). Focusing on the peak at $M\sim 67$, the results shows a slightly better performance (lower cost) with $L=1$ for the constant rewiring case, while negligible differences when $L=3$. Figures 8(b) and 8(c) show these same results for different average degrees $\mu$, where coevolution itself does not work. One can see on these results that constant rewiring does not work either.
Figure 8: Comparison between coevolution and constant rewiring dynamics. Computational cost $\left\langle C\right\rangle$ as a function of network size $M$ for different values of the average degree. The green line is the IS (Eq. 4), the black dashed line is the FCN result and RC is the reference case (equivalent to $L=0$). (a) Scenarion where coevolution works: average degree $\mu=10$. Scenarios where coevolution does not work: (b) $\mu=5$ and (c) $\mu=15$.
Lastly we studied the impact of non-linearity on the computational cost, which happens when $q$ is different from 0 or 1 [38]. In this work non-linearity yield no significant outcome, as can be shown on Fig. 9. The case $q=0.5$ only showed slightly better performance than the linear case $q=1$ around $M=67$. Other than this region there were no significant differences.
Figure 9: Study of the impact of non-linear effects on the coevolution dynamics. Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of the non-linear parameter $q$. Parameters: $\mu=10$, $S=3000$ and $L=3$. The violet circles correspond to $q=0.5$, the orange circles to $q=1.5$, and the red circles to $q=2$. The reference case RC ($L=0$ and $q=1$) is depicted in gray. The non-linear case refers to $q\neq 0.0$ and $q\neq 1.0$.
@@ 4 Conclusions
We investigated the impact of coevolution on a cooperative process, where the target can substitute a few of his neighbors in order to solve the problem more rapidly. The neighbors are important because copy one of them (exploitation) is one possible strategy to improve the fitness. So, each agent analyses the fitness of his neighbors and uses this piece of information to decide wether or not to substitute a few of them. The number of neighbors to be replaced is an input parameter of the model. This creates a tunable mechanism enabling some sort of intensity on the coevolution process: a “weak” coevolution would be $L=1$, which has been studied on different social models. A “stronger” coevolution would be $L\geq 2$, which, to the best of our knowledge, has not been studied before.
A natural strategy for the target is to change some of his neighbors if they have worst (smaller) fitness than him. Nonetheless, this rule did not give satisfactory results in our tests when compared with the reference case. Surprisingly, the opposite rule did yield some interesting results. We chose to apply this new coevolution model on a cooperative process. Using as prototype the NK model, our results showed a convoluted behavior of the system efficiency.
At the easy landscape, coevolution did bring significantly better results only at lower average degree (Fig. 4(a)). Additionally, the larger $L$ the lower the $C$, except at $L=1$ at which the results are worse. However, the FCN case is still the more efficient one at $K=0$, regardless other parameters. At the difficult landscape $K=4$, the behaviour was unanticipated. At lower average degree, after a cost increase at $L=1$, the cost cutback for increasing $L$ for most part of the system size range. Except at around $M\sim 70$ where a peak starts to grow at $L=5$ and becomes higher than the IS cost ((Fig. 4(b)). At higher average degree, this peak on the cost already exists at $L=0$ (Fig. 5(b)). It disappears at $L=3$ and comes back at $L=5$. Indeed, this minimum can be better visualized on the $(M,L)$ space (Fig. 6(a)). Fixing $M=67$, the minimum of the computational cost can be seen in details on the 2D map of Fig. 7(b) (blue region).
This work showed that this tunable coevolution can drastically change the behavior of computational cost. It opens some interesting perspectives such as the impact of this tunable coevolution on the phase diagram involving absorbing phase transition of models such as Axelrod’s, voter model and others.
@@ Acknowledgements
The model was implemented using Fortran and the analysis and graphics were developed using Python (packages: Numpy [50], Matplotlib [51], NetworkX [52] and Pandas [53]) and Gnuplot.
Appendix A Appendix
A.1 NK landscape
In this appendix we present a simpled description of how the function $f(\mathcal{X})$ is created for a given landscape.
Following the formal definition from Eq. 1, we have $f:\mathscr{D}\rightarrow\mathscr{R}$, where $\mathscr{D}$ is the set of $2^{N}$ binary vectors $\mathcal{X}$ and $\mathscr{R}$ is the real numbers set. The argument of the contributing functions $\Phi_{i}$ is given by Eq. 2. To better understand this, suppose the vector $\mathcal{Y}=(y_{0},y_{1},y_{2},y_{3})=(0,1,1,0)$ on the landscape $N=4$ and $K=2$. There will be $K+1=3$ bits on the argument of each $\Phi_{i}$. The first bit from the i-th contribution is the i-th bit of $\mathcal{Y}$, while the other bits are the $K=2$ neighbors on the right. So, the first contribution of $\mathcal{Y}$ is $\Phi_{0}(y_{0},y_{1},y_{2})=(0,1,1)$. The second contribution will be $\Phi_{1}(y_{1},y_{2},y_{3})=\Phi_{1}(1,1,0)$. On the third contribution $\Phi_{2}$, the third bit will be the fifth one of $\mathcal{Y}$. However, and the string $\mathcal{Y}$ has only $N=4$ bits, so the fifth bit will be the first one. This equivalent to subtract $N$ from the position of the bit: 5-4 = 1. Thus: $\Phi_{2}(y_{2},y_{3},y_{0})=\Phi_{2}(1,0,0)$. In the same way: $\Phi_{3}(y_{3},y_{0},y_{1})=\Phi_{2}(0,0,1)$ Putting together the four contributions, we have:
$f(0,1,1,0)$ $=$ $\tfrac{1}{4}\Phi_{0}(0,1,1)+\tfrac{1}{4}\Phi_{1}(1,1,0)$ (9)
$+$ $\tfrac{1}{4}\Phi_{2}(1,0,0)+\tfrac{1}{4}\Phi_{3}(0,0,1).$
Other examples are:
$f(0,0,0,0)$ $=$ $\tfrac{1}{4}\Phi_{0}(0,0,0)+\tfrac{1}{4}\Phi_{1}(0,0,0)$
$+$ $\tfrac{1}{4}\Phi_{2}(0,0,0)+\tfrac{1}{4}\Phi_{3}(0,0,0),$
$f(0,0,1,0)$ $=$ $\tfrac{1}{4}\Phi_{0}(0,0,1)+\tfrac{1}{4}\Phi_{1}(0,1,0)$
$+$ $\tfrac{1}{4}\Phi_{2}(1,0,0)+\tfrac{1}{4}\Phi_{3}(0,0,0),$
$f(1,0,1,0)$ $=$ $\tfrac{1}{4}\Phi_{0}(1,0,1)+\tfrac{1}{4}\Phi_{1}(0,1,0)$
$+$ $\tfrac{1}{4}\Phi_{2}(1,0,1)+\tfrac{1}{4}\Phi_{3}(0,1,0),$
$f(0,1,1,1)$ $=$ $\tfrac{1}{4}\Phi_{0}(0,1,1)+\tfrac{1}{4}\Phi_{1}(1,1,1)$
$+$ $\tfrac{1}{4}\Phi_{2}(1,1,0)+\tfrac{1}{4}\Phi_{3}(1,0,1),$
$f(1,1,1,1)$ $=$ $\tfrac{1}{4}\Phi_{0}(1,1,1)+\tfrac{1}{4}\Phi_{1}(1,1,1)$
$+$ $\tfrac{1}{4}\Phi_{2}(1,1,1)+\tfrac{1}{4}\Phi_{3}(1,1,1).$
It is worth mention that each argument on the functions $\Phi_{i}$ contains $K+1$ bits. So, there are $2^{K+1}$ possible combinations for the $N$ contributions from each vector.
A.2 Base 2 and 10
Conceptually, the description from previous section is enough. However, computationally, it is compelling to represent the vectors $\mathcal{X}$ and the arguments of the contributions $\Phi_{i}$ as integers. Both of them are binary sequences, 0s and 1s, numbers in base 2, and thus, can be converted to decimal numbers (base 10). In general, considering $N=4$: $(x_{0},x_{1},x_{2},x_{3})=x_{3}\cdot 2^{0}+x_{2}\cdot 2^{1}+x_{1}\cdot 2^{2}+x_{0}\cdot 2^{3}$, starting the count from the right. Thus,
$(0,0,0,0)=0,\qquad(0,0,0,1)=1,$
$(0,0,1,0)=2,\qquad(0,1,0,0)=4,$
$(1,0,0,0)=8,\qquad(0,1,0,1)=5.$
Computationally, this is much more efficient because we can record each string by its base 10 correspondent number (an integer), which takes much less memory than the string with $N$ bits.
We can do the same for the arguments of the contribution functions. Eq. 9 would be:
$f(0,1,1,0)$ $=$ $f(6),$
$=$ $\frac{1}{4}\left[\Phi_{0}(3)+\Phi_{1}(6)+\Phi_{2}(4)+\Phi_{3}(1)\right].$
Other example: the fitness of the vector $\mathcal{X}=7$ considering $N=4$ and $K=3$ is:
$f(7)$ $=$ $f(0,1,1,1),$
$=$ $\tfrac{1}{4}\Phi_{0}(0,1,1)+\tfrac{1}{4}\Phi_{1}(1,1,1)$
$+$ $\tfrac{1}{4}\Phi_{2}(1,1,0)+\tfrac{1}{4}\Phi_{3}(1,0,1),$
$=$ $\frac{1}{4}\left[\Phi_{0}(3)+\Phi_{1}(7)+\Phi_{2}(6)+\Phi_{3}(5)\right].$
This binary to decimal conversion is performed only once, when the landscaped is built, in the beginning of the algorithm. The fitness is the function $f(i)$ where $i=0,1,2,3,...,2^{N}-1$. Likewise, the contributions are functions $\Phi_{i}(j)$ where $i=0,1,2,3,...,N-1$ and $j=0,1,2,3,...,2^{K+1}-1$. Initially, the vectors $\mathcal{X}$ are randomly generated, which can be done by generating random integers from 0 to $12^{N}-$ 1 using a uniform distribution.
Once the arguments are defined, we have to define the functions $\Phi_{i}(j)$ itself. In this work, we use the NK landscape for a cooperative problem where the simulation comes to a halt when one agent finds the global maximum. So, it is interesting to keep the fitness on the range $\left[0,1\right]$. Thus, we define the functions $\Phi_{i}(j)$ as a random number from the uniform distribution on the same range $\left[0,1\right]$.